### v3

@@ 6 Tidal asteroseismology: the interplay between rotation, tides, and oscillations in binaries
The classes of observed stars exhibiting tidally excited and tidally perturbed oscillations open up the new window of tidal asteroseismology, irrespective of the rotation rate, from ultra-slow to very fast. Progress in the forward modelling of the internal physical properties from carefully identified modes is the next yet extremely challenging step, given the numerous orbital parameters involved, as well as the numerical aspects involved in modal decompositions and the computation of damping coefficients, and the complex physical phenomenon of pseudo-synchronisation (Townsend & Sun 2023). While a concerted plan-of-attack for tidal asteroseismic modelling is still awaited, applications to large samples of close binaries are the best way to understand and calibrate tidal dissipation and close binary evolution. Achieving this would be a major step ahead as input for proper binary population synthesis and gravitational-wave progenitor modelling (cf. Fig. 1). The modern public tool to model stellar tides developed by Sun et al. (2023a) is highly relevant in this context.
7 The special cases of subdwarf and merger seismology
@@
Subdwarfs play a special role in terms of close binarity, as their bare existence demands binary interaction at evolved stages of stellar evolution, including mergers (Han et al. 2003). Indeed, either they occur in a binary or they are single subdwarfs as the result of a merger (Heber 2009). Just as is the case for white dwarfs, asteroseismology of subdwarfs has advanced appreciably thanks to short-cadence (1 min and 2 min or 20 sec) light curves assembles with Kepler, K2, and TESS (Hermes et al. 2017; Bell et al. 2019; Reed et al. 2021; Uzundag et al. 2021; Córsico et al. 2022; Uzundag et al. 2022; Romero et al. 2022), including post-merger subdwarfs (Vos et al. 2021). Asteroseismic probing of almost all subwarfs and all white dwarfs is done from modelling of their identified g modes via period spacing patterns and/or rotationally split triplets and quintuplets, where the Coriolis and tidal forces can be treated as a small perturbation for the computation of the modes because the periods of all detected modes (minutes to hours) are much shorter than the rotation or orbital periods (days).
@@
A review of white dwarf asteroseismology was presented by Córsico et al. (2019), but no such comprehensive summary is available for subdwarfs. This is far beyond the scope of this paper, hence we just highlight some recent findings for subdwarf binaries. Two major conclusions connected with transport processes can be drawn from published subdwarf asteroseismology:
@@
the angular momentum of the convective core of red giants and of subdwarfs is in agreement with the angular momentum of white dwarfs (Aerts et al. 2019);
@@
subdwarfs and white dwarfs have respectively more massive helium and carbon-oxygen cores compared to those predicted by standard stellar evolution models without extra mixing beyond the Schwarzschild boundary. Hence core boundary mixing is prominent throughout the evolution of low- and intermediate-mass stars (Van Grootel et al. 2010a, b; Charpinet et al. 2011; Hermes et al. 2017; Charpinet et al. 2019b, a; Giammichele et al. 2018; Uzundag et al. 2021; Giammichele et al. 2022, where the latter review paper places both types of compact pulsators and the properties of their carbon-oxygen cores into an evolutionary context).
@@
Future progress in the calibration of close binary evolution at evolved phases can come from several aspects. First of all, many of the modes of subdwarfs and white dwarfs exhibit amplitude and frequency modulations on timescales of weeks to months, pointing to non-linear resonant mode interactions. The exploitation of such interactions is easier than for the rapidly rotating SPB and Dor stars discussed in Sect. 4.4 and is therefore somewhat more advanced (Zong et al. 2016b, a). Yet, much remains to be done in the development and application of non-linear asteroseismology, also for the compact pulsators. The recent finding of differential rotation in the subdwarf EPIC 220422705 by Ma et al. (2022) is an important step towards future modelling of its non-linear modes.
@@
As for the calibration of binary evolution theory, several recent key findings were delivered by asteroseismology. First of all, short-period subdwarf–M dwarf binaries are not necessarily synchronised. A key laboratory for further study is TIC 137608661, whose internal rotation period of 4.6 d deduced from asteroseismology is much longer than the short orbital period of only about 0.3 d. The asteroseismic study of this subdwarf binary was placed into context of subdwarf rotation and binary properties by Silvotti et al. (2022). Their Table 4 and Fig. 16 suggests differential rotation for most of the subdwarfs in short-period binaries with either an M-dwarf or a white dwarf companion. Another striking result comes from the tidally tilted oscillations (cf. Fuller et al. 2020, as discussed in the previous section) of the binary subdwarf HD 265435 discovered and analysed by Jayaraman et al. (2022). Its white dwarf companion revolves around it in only 1.65 hr. The subdwarf was found to be near or just beyond the end of its central helium burning, hence this object is a Type Ia supernova progenitor. Most – but not all – of the subdwarf’s oscillations have frequencies that are exact multiples of the orbital frequency. This star was diagnosed with acoustic modes having periods around 250 s. So far these tidally tilted oscillations were modelled while ignoring the Coriolis and centrifugal forces. However, the tidal distortion for a binary with such a tight orbit shifts the mode frequencies appreciably compared to the case of a spherical star as the authors have assumed. Future measurements of the internal and surface rotation for more of such short-period binaries are key to guide theories of tidal synchronisation.
@@
Backtracking to earlier phases, Li et al. (2022c) used asteroseismology of red clump red giants to identify two groups of stripped stars due to binary interactions. The first group comprises several underluminous red giants with smaller than usual helium burning cores, which are interpreted as resulting from heavy mass loss due to envelope stripping by a companion while they ascended the red giant branch. The second group consists of tens of red giants whose ages would exceed the age of the Universe if interpreted from single star evolution with no or nominal mass loss on the red giant branch. The small sample sizes of these two groups are consistent with binary fractions as in Fig. 2. Coupling the properties of these red giants to models of post-mass-transfer binary systems will be invaluable to calibrate binary evolution theories of low- to intermediate-mass stars.
@@
Moving on to the even earlier phases and to more massive stars, Fig. 2 and binary evolution calculations by Schneider et al. (2020) tell us that 20% of OB stars come into contact with their companion. Hence a large population of binary mergers must exist, either formed from collision or from more gentle coalescence (Schneider et al. 2019). The properties of such mergers are highly uncertain and searching for such products is difficult. Asteroseismology could play a major role here, both in terms of establishing a proper calibration for the statistics of the sample of high-mass mergers and of probing their internal structure. This would allow to check if mergers indeed have a strong large-scale magnetic field as the MHD simulations by Schneider et al. (2019) predict. The first calculations of seismic diagnostics for stellar mergers were compared with those of single blue supergiants in the hydrogen shell burning phase (so-called Hertzsprung Gap (HG) in the HRD), of main-sequence stars with an exceptionally high level of core overshooting, and of evolved single blue giants undergoing a blue loop during their core helium burning by Bellinger et al. (2024). These authors computed grids of non-rotating equilibrium models and their oscillation frequencies for these four different populations, considering the mass range M. This theoretical study shows that asteroseismology in principle offers the potential to distinguish among these four cases from the large frequency separation of identified p modes and the average period spacing of identified g modes. A major point to be improved is to include internal rotation in such models.
@@
To illustrate the potential of “merger seismology” (a term suggested by Dr. Fabian Schneider from the Heidelberg Institute for Theoretical Studies) from a practical point of view, we show in Fig. 9 predicted seismic g-mode period spacings of a merger product in the HG and compare them to those of a single blue giant residing in the same part of the HRD (Henneco et al., in prep.). It can be seen that the period spacing values of these two models are clearly different in the sense that the average period spacing of the merger product is lower than the one of a model resulting from nominal single star evolution. This is in agreement with the results found by Bellinger et al. (2024). Moreover, the oscillations of the merger shown in Fig. 9 lead to different mode periodicities and dips structures in the period spacing patterns compared with those of the single HG star. The shorter mode periods and the sharper dips at high radial order for the merger are due to particular structures in its Brunt-Väisälä frequency, originating from the chemical properties near its core reflected in its gradient. The occurrence or absence of intermediate convection zones, often caused by semi-convection, also plays a role. The profile of for the two models with similar position in the HRD are markedly different, resulting in different mode trapping. Signatures as in Fig. 9, when computed for a representative population of mergers with proper inclusion of internal rotation, will provide a powerful tool to search for merger products in a machine-learning context when applied to Kepler and TESS light curves. Once found, modelling of the morphology of the period spacing pattern(s) would allow us to assess the profiles of the internal chemical composition, rotation, and magnetic field of mergers.
@@
TESS is currently monitoring about 1500 massive OB dwarfs, giants, and supergiants across the sky at a cadence of 2 min. Figure 2 makes it clear that dedicated searches for signals of merger products among these stars based on predictions as in Fig. 9 offer an interesting route for massive star asteroseismology. Future searches in the TESS data are highly relevant to calibrate populations of gravitational wave progenitors seismically because mergers should also occur in what were originally doubles, triples or multiple systems in general. The discovery of 78 Cep stars in eclipsing binaries, including 59 new ones, by Eze & Handler (submitted) offer a fantastic new data set to apply the principles in Fig. 8 and to predict how many of these targets will evolve into a merger, before or after the supernova explosion(s) that will be happening during their future evolution. Discovering mergers from seismology, unravelling their internal structure, and comparing it to the structure of massive SPB or Cep detached or semi-detached binaries offers a great way to calibrate tidal dissipation and close binary evolution. It would also provide an elegant and independent way to evaluate the measured mass distribution of the black hole and neutron star mergers that lead to gravitational wave signals (cf. Fig. 1).
@@ 8 Onward to a bright future for asteroseismology, for slow and fast rotators
The TESS mission keeps on going and provides ever longer time-series photometry covering many years by now (albeit with interruptions of a year for each of its covered sectors). Moreover, the launch of the PLATO mission is on the horizon (end 2026 at the time of writing). With its core program focused on exoplanet host candidates and its open Guest Observer program, it will cover all of slow to fast rotators in all evolutionary phases, from single stars through binaries and multiples to clusters of various ages and metallicities (Rauer et al., submitted). Even without the additional candidate space missions based on a similar principle of high-cadence high-precision uninterrupted space photometry, such as HAYDN (Miglio et al. 2021) and ET (Ge et al. 2022), TESS and PLATO guarantee a bright future for asteroseismology for many years to come. A wealth of suitable light curves covering many years with high duty cycle will become available for millions of stars. These data feeds the fountain of opportunities shown in Fig. 1 and will allow us to push asteroseismology to high-mass stars (unavoidably in (close) binary systems, cf. Fig. 2), covering from early to evolved evolutionary phases, with some on their way to become gravitational wave sources.
@@
As we have tried to stress in this review, asteroseismology based on numerous low-frequency modes in moderate to fast rotators requires long and uninterrupted data strings to reach sufficient frequency precision and to unravel long and complex beating patterns (cf. Fig. 7). Even the 4-yr uninterrupted Kepler light curves are insufficient to resolve some of the individual modes of Dor, SPB, and Be pulsators. Years-long time strings are also necessary to detect and model the internal rotation and magnetic field profiles inside stars.
@@
8.1 Pushing Gaia beyond its limits to assist in asteroseismology of intermediate- and high-mass stars
@@
The Gaia space mission of the European Space Agency (Gaia Collaboration et al. 2016) is not made to do asteroseismology based on non-radial oscillations. Indeed, high sampling rates and continuous monitoring with duty cycles above 90% are required to push the Nyquist frequency of the light curves to high enough values and to identify a sufficient number of oscillation modes. Gaia cannot deliver that, nor can it offer mode detections for amplitudes below mmag as required to detect solar-like oscillations. Nevertheless, Gaia has a significant role to play, particularly for asteroseismology of moderate and fast rotators. All asteroseismic modelling efforts nowadays benefit from Gaia’s data, notably the high-precision parallaxes (Gaia Collaboration et al. 2018) leading to luminosities that allow us to limit the computation of grids of equilibrium models required for the modelling. For slowly rotating low-mass stars, the Gaia effective temperatures additionally lead to fairly precise radius estimates, which can be confronted with the asteroseismic ones. Conversely, combined asteroseismic radii and Gaia effective temperatures lead to asteroseismic distance estimates of remarkable precision, up to distances of kilo-parsecs (Huber et al. 2017; Davies et al. 2017).
The continued improvement of the Gaia, Kepler and TESS data reductions provide excellent means for more accurate instrument calibrations, with ever better future capacity for large and homogeneous population studies. Along with the Gaia distances and proper motions, precise asteroseismic ages are particularly valuable for galactic archaeology studies (cf. Fig. 1). In addition, we also point to other aspects of the future role that Gaia can play, by pushing beyond nominal use of its data. The three light curves made available in Gaia’s Data Release 3 (DR3) in the G, RP, and BP photometric bands delivered millions of new variable stars (Eyer et al. 2023), among which thousands of candidates with non-radial oscillations having amplitudes above a few mmag (Gaia Collaboration et al. 2023a). Figure 10 shows the position of more than 25000 new non-radial pulsators in the upper HRD discovered from Gaia DR3. The more than 11000 new g-mode pulsators among these have properties fully in line with the few hundreds of the Kepler pulsators in these classes (Aerts et al. 2023) and thus offer great potential for future ‘industrialised’ ensemble asteroseismology of intermediate- and high-mass stars. The requirement is that sufficient non-radial modes can be identified from dedicated space photometry by Kepler, TESS or PLATO, in addition to their dominant mode(s) discovered in the Gaia photometry. This is already the case for about 700 moderate and fast rotators observed by Kepler (Aerts 2021) and will be extended to thousands of such new non-radial pulsators currently under scrutiny by TESS (Pedersen et al. 2019; Antoci et al. 2019; Cunha et al. 2019; Burssens et al. 2020; Skarka et al. 2022; Garcia et al. 2022b, and Hey & Aerts, in prep.).
@@
Fig. 10 includes a comparison with the positions of four instability strips where non-radial modes are expected to be excited based on one particular choice of input physics for each of the Cep, SPB, Sct, and Dor pulsator classes. We know nowadays from asteroseismology of moderate and fast rotators that such comparisons with one strip per class based on only one choice of input physics is not that meaningful, in view of the large variety of internal rotation rates and mixing levels in the interior of single and binary stars in that part of the HRD (e.g., Moravveji et al. 2015, 2016; Szewczuk & Daszyńska-Daszkiewicz 2018; Wu & Li 2019; Wu et al. 2020; Mombarg et al. 2020; Pedersen et al. 2021; Pedersen 2022b; Szewczuk et al. 2022). That is why many stars occur outside of the particular strip chosen to illustrate the comparison in Fig. 10. Any good comparison between data and theory should include the diversity of input physics that got revealed by asteroseismology, notably the various choices for the adopted opacity tables, for chemical mixtures and metallicities, for the mixing length parameter of time-dependent convection calculations, and keeping in mind moderate and fast rotation (Townsend 2005; Dupret et al. 2005c, a, b; Moravveji 2016; Szewczuk & Daszyńska-Daszkiewicz 2017, to list just a few papers).
Following Fig. 2, a large fraction of the stars inside and in between the instability strips shown in Fig. 10 must be binaries. Their deduced luminosity and effective temperature may thus be inaccurate if the binarity remains unknown. Gaia’s capability to find binarity from HRDs or CMDs will trigger follow-up spectroscopy, while dedicated high-cadence photometry will allow us to validate the seismic activity. Once the binarity is confirmed (or refuted) it will help delineate the input physics of stellar models for single and binary stars within the four classes of pulsators shown in Fig. 10. This can be achieved by demanding that the overall areas covered by an assembly of instability strips computed for a variety of opacity tables, metallicities, and transport processes delivered by asteroseismology encapsulate all the detected and confirmed single and binary pulsators.

### v4

@@ 6 Tidal asteroseismology: Interplay between rotation, tides, and oscillations in binaries
The classes of observed stars exhibiting tidally excited and tidally perturbed oscillations open up the new window of tidal asteroseismology, irrespective of the rotation rate, from ultra-slow to very fast. Progress in the forward modelling of the internal physical properties from carefully identified modes is the next extremely challenging step, given the numerous orbital parameters involved, as well as the numerical aspects involved in modal decompositions and the computation of damping coefficients, and the complex physical phenomenon of pseudo-synchronisation (Townsend & Sun 2023). While we are still waiting for a concerted plan of attack for tidal asteroseismic modelling, applications to large samples of close binaries are the best way to understand and calibrate tidal dissipation and close binary evolution. Achieving this would be a major step ahead as input for proper binary population synthesis and gravitational-wave progenitor modelling (see Fig. 1). The modern public tool to model stellar tides developed by Sun et al. (2023a) is highly relevant in this context.
7 The special cases of sub-dwarf and merger seismology
@@
Sub-dwarfs play a special role in terms of close binarity as their bare existence demands binary interaction at evolved stages of stellar evolution, including mergers (Han et al. 2003). Either they occur in a binary or they are single sub-dwarfs as the result of a merger (Heber 2009). Just as is the case for white dwarfs, the asteroseismology of sub-dwarfs has advanced appreciably thanks to short-cadence (1 min, 2 min, or 20 sec) light curves assembled with Kepler, K2, and TESS (Hermes et al. 2017; Bell et al. 2019; Reed et al. 2021; Uzundag et al. 2021; Córsico et al. 2022; Uzundag et al. 2022; Romero et al. 2022), including post-merger sub-dwarfs (Vos et al. 2021). Asteroseismic probing of almost all sub-dwarfs and all white dwarfs is done from the modelling of their identified g modes via period spacing patterns and/or rotationally split triplets and quintuplets, where the Coriolis and tidal forces can be treated as a small perturbation for the computation of the modes because the periods of all detected modes (minutes to hours) are much shorter than the rotation or orbital periods (days).
@@
A review of white dwarf asteroseismology was presented by Córsico et al. (2019), but no such comprehensive summary is available for sub-dwarfs. This is far beyond the scope of this paper, and hence we just highlight some recent findings for sub-dwarf binaries. Two major conclusions connected with transport processes can be drawn from published sub-dwarf asteroseismology:
@@
The angular momentum of the convective core of red giants and of sub-dwarfs is in agreement with the angular momentum of white dwarfs (Aerts et al. 2019);
@@
Sub-dwarfs and white dwarfs have respectively more massive helium and carbon–oxygen cores compared to those predicted by standard stellar evolution models without extra mixing beyond the Schwarzschild boundary. Hence, core boundary mixing is prominent throughout the evolution of low- and intermediate-mass stars (Van Grootel et al. 2010a, b; Charpinet et al. 2011; Hermes et al. 2017; Charpinet et al. 2019b, a; Giammichele et al. 2018; Uzundag et al. 2021; Giammichele et al. 2022, where the latter review paper places both types of compact pulsators and the properties of their carbon-oxygen cores into an evolutionary context).
@@
Future progress in the calibration of close binary evolution at evolved phases can come from several aspects. First of all, many of the modes of sub-dwarfs and white dwarfs exhibit amplitude and frequency modulations on timescales of weeks to months, pointing to non-linear resonant mode interactions. The exploitation of such interactions is easier than for the rapidly rotating SPB and Dor stars discussed in Sect. 4.4, and is therefore somewhat more advanced (Zong et al. 2016b, a). However, much remains to be done in the development and application of non-linear asteroseismology, also for the compact pulsators. The recent finding of radial-differential rotation between the core and surface at the level of about 1.2 in the sub-dwarf EPIC 220422705 by Ma et al. (2022) is an important step towards the future modelling of its non-linear modes.
@@
As for the calibration of binary evolution theory, several recent key findings were delivered by asteroseismology. First of all, short-period sub-dwarf–M dwarf binaries are not necessarily synchronised. A key laboratory for further study is TIC 137608661, whose core rotates slower than its surface, with a period of 4.6 d deduced from g-mode asteroseismology. This core rotation period is much longer than the short orbital period of only about 0.3 d. The asteroseismic study of this sub-dwarf binary was placed into the context of sub-dwarf rotation and binary properties by Silvotti et al. (2022). Their Table 4 and Fig. 16 suggests close to rigid rotation for most of the other sub-dwarfs in short-period binaries with either an M-dwarf or a white dwarf companion. Another striking result comes from the tidally tilted oscillations (see Fuller et al. 2020, as discussed in the previous section) of the binary sub-dwarf HD 265435 discovered and analysed by Jayaraman et al. (2022). Its white dwarf companion revolves around it in only 1.65 hr. The sub-dwarf was found to be near or just beyond the end of its central helium burning, and hence this object is a Type Ia supernova progenitor. Most, but not all, of the sub-dwarf’s oscillations have frequencies that are exact multiples of the orbital frequency. This star was diagnosed with acoustic modes having periods of around 250 s. To date, these tidally tilted oscillations have been modelled, while the Coriolis and centrifugal forces have been ignored. However, the tidal distortion for a binary with such a tight orbit shifts the mode frequencies appreciably compared to the case of a spherical star as the authors have assumed. Future measurements of the internal and surface rotation for more short-period binaries are key to guiding theories of tidal synchronisation.
@@
Backtracking to earlier phases, Deheuvels et al. (2022), Li et al. (2022c), and Rui & Fuller (2024) used asteroseismology of red clump red giants to identify groups of stripped stars due to binary interactions. One group comprises several underluminous red giants with smaller than usual helium burning cores, which are interpreted as resulting from heavy mass loss due to envelope stripping by a companion while they ascended the red giant branch. Another group consists of tens of red giants whose ages would exceed the age of the Universe if interpreted from single-star evolution with no or nominal mass loss on the red giant branch. However, another mechanism leads to core-helium burning red giants with higher than usual luminosity as a result of the consumption of a white dwarf during the earlier main-sequence phase. The small sample sizes of these observed anomalous groups are consistent with binary fractions as in Fig. 2. Coupling the properties of these red giants to models of post-mass-transfer binary systems will be invaluable to calibrate binary evolution theories of low- to intermediate-mass stars.
@@
Moving on to the even earlier phases and to more massive stars, Fig. 2 and binary evolution calculations by Schneider et al. (2020) tell us that 20% of OB stars come into contact with their companion. Hence, a large population of binary mergers must exist, either formed from collision or from more gentle coalescence (Schneider et al. 2019). The properties of such mergers are highly uncertain and searching for such products is difficult. Asteroseismology could play a major role here, both in terms of establishing a proper calibration for the statistics of the sample of high-mass mergers and of probing their internal structure. This would allow us to check whether mergers indeed have a strong large-scale magnetic field, as the MHD simulations by Schneider et al. (2019) predict. The first calculations of seismic diagnostics for stellar mergers were compared with those of single blue supergiants in the hydrogen shell burning phase (called the Hertzsprung gap (HG) in the HRD), of main-sequence stars with an exceptionally high level of core overshooting, and of evolved single blue giants undergoing a blue loop during their core-helium burning by Bellinger et al. (2024). These authors computed grids of non-rotating equilibrium models and their oscillation frequencies for these four different populations, considering the mass range M⊙. This theoretical study shows that asteroseismology in principle offers the potential to distinguish among these four cases from the large frequency separation of identified p modes and the average period spacing of identified g modes. A major point to be improved is to include internal rotation in such models.
@@
To illustrate the potential of ‘merger seismology’ (a term suggested by Dr. Fabian Schneider from the Heidelberg Institute for Theoretical Studies) from a practical point of view, we show in Fig. 9 the predicted seismic g-mode period spacings of a merger product in the HG and compare them to those of a single blue giant residing in the same part of the HRD, following the study by Henneco et al. (2024). It can be seen that the period spacing values of these two models are clearly different in the sense that the average period spacing of the merger product is lower than that of a model resulting from nominal single-star evolution. This is in agreement with the results found by Bellinger et al. (2024). Moreover, the oscillations of the merger shown in Fig. 9 lead to different mode periodicities and dip structures in the period spacing patterns compared with those of the single HG star. The shorter mode periods and the sharper dips at high radial order for the merger are due to particular structures in its Brunt-Väisälä frequency, originating from the chemical properties near its core reflected in its -gradient. The occurrence or absence of intermediate convection zones, often caused by semi-convection, also plays a role. The profile of for the two models with similar positions in the HRD are markedly different, resulting in different mode trapping and mode coupling according to the occurrence of various mode cavities. Signatures such as those in Fig. 9, when computed for a representative population of mergers with proper inclusion of internal rotation, will provide a powerful tool to search for merger products in a machine-learning context when applied to Kepler and TESS light curves. Once found, modelling the morphology of the period spacing pattern(s) would allow us to assess the profiles of the internal chemical composition, rotation, and magnetic field of mergers.
@@
TESS is currently monitoring about 1500 massive OB dwarfs, giants, and supergiants across the sky at a cadence of 2 min. Figure 2 makes it clear that dedicated searches for signals of merger products among these stars based on the predictions in Fig. 9 offer an interesting route for massive star asteroseismology. Future searches in the TESS data are highly relevant to calibrating populations of gravitational wave progenitors seismically because mergers should also occur in what were originally double, triple, or multiple systems in general. The discovery of 78 Cep stars in eclipsing binaries, including 59 new pulsators, by Eze & Handler (2024) offers a fantastic new data set to apply the principles in Fig. 8 and to predict how many of these targets will evolve into a merger before or after the supernova explosion(s) that will be happening during their future evolution. Discovering mergers from seismology, unravelling their internal structure, and comparing it to the structure of massive SPB or Cep detached or semi-detached binaries offers a great way to calibrate tidal dissipation and close binary evolution. It would also provide an elegant and independent way to evaluate the measured mass distribution of the black hole and neutron star mergers that lead to gravitational wave signals (see Fig. 1).
@@ 8 Onward to a bright future for asteroseismology, for slow and fast rotators
The TESS mission keeps on going, and currently provides ever longer time series photometry covering many years (albeit with interruptions of a year for each of its covered sectors). Moreover, the launch of the PLATO mission is on the horizon (late 2026, at the time of writing). With its core programme focused on exoplanet host candidates and its open Guest Observer programme, it will cover slow to fast rotators in all evolutionary phases, from single stars through binaries and multiples to clusters of various ages and metallicities (Rauer et al. 2024). Even without the additional candidate space missions based on a similar principle of high-cadence high-precision uninterrupted space photometry, such as the High-precision AsteroseismologY of DeNse stellar fields mission (HAYDN, Miglio et al. 2021) and the mission ExtraTerrestrial: To Find the First Earth 2.0 (ET, Ge et al. 2022), TESS and PLATO guarantee a bright future for asteroseismology for many years to come. A wealth of suitable light curves covering many years with high duty cycle will become available for millions of stars. These data feed the fountain of opportunities shown in Fig. 1 and will allow us to push asteroseismology to high-mass stars (unavoidably in close binary systems; see Fig. 2), covering from early to evolved evolutionary phases, with some on their way to becoming gravitational wave sources.
@@
As we have tried to emphasise in this review, asteroseismology based on numerous low-frequency modes in moderate to fast rotators requires long and uninterrupted data strings to reach sufficient frequency precision and to unravel long and complex beating patterns (see Fig. 7). Even the 4 yr uninterrupted Kepler light curves are insufficient to resolve some of the individual modes of Dor, SPB, and Be pulsators. Years-long time strings are also necessary to detect and model the internal rotation and magnetic field profiles inside stars.
@@
8.1 Pushing Gaia beyond its limits to assist in the asteroseismology of intermediate- and high-mass stars
@@
The Gaia space mission of the European Space Agency (Gaia Collaboration et al. 2016) was not designed to do asteroseismology. High sampling rates and continuous monitoring with duty cycles above 90% are required to perform asteroseismology. This is needed to push the Nyquist frequency of the light curves to high enough values, and to detect and identify a sufficient number of oscillation modes. Gaia cannot deliver that. In particular it cannot offer mode detections for amplitudes below the millimagnitude range as required to detect solar-like oscillations. Nevertheless, Gaia has a significant role to play, particularly for the asteroseismology of moderate and fast rotators.
All asteroseismic modelling efforts currently benefit from Gaia’s data, notably the high-precision parallaxes (Gaia Collaboration et al. 2018), leading to luminosities that guide the computation of grids of equilibrium models required for the modelling. For slow-rotating low-mass stars, the Gaia effective temperatures additionally lead to fairly precise radius estimates, which can be confronted with the asteroseismic values. Conversely, combined asteroseismic radii and Gaia effective temperatures lead to asteroseismic distance estimates of remarkable precision, up to distances of kiloparsecs (Huber et al. 2017; Davies et al. 2017).
The continued improvement of the Gaia, Kepler and TESS data reductions provide excellent means for more accurate instrument calibrations, with ever better future capacity for large and homogeneous population studies. Along with the Gaia distances and proper motions, precise asteroseismic ages are particularly valuable for galactic archaeology studies (see Fig. 1).
@@
We also point to additional roles that Gaia can play by pushing beyond the nominal use of its data. The three light curves made available in Gaia’s Data Release 3 (DR3) in the G, RP, and BP photometric bands delivered millions of new variable stars (Eyer et al. 2023), among which thousands of candidates with non-radial oscillations having amplitudes above a few mmag (Gaia Collaboration et al. 2023a). Figure 10 shows the position of more than 100,000 new (non-)radial pulsators in the upper HRD discovered from Gaia DR3. This figure includes a comparison with the positions of four instability strips where non-radial modes are expected to be excited based on one particular choice of input physics for each of the Cep, SPB, Sct, and Dor pulsator classes. We know from asteroseismology that such comparisons with one strip per class based on only one choice of input physics is too limited, in view of the large variety of internal rotation rates and mixing levels in the interior of single and binary stars in that part of the HRD (e.g. Moravveji et al. 2015, 2016; Szewczuk & Daszyńska-Daszkiewicz 2018; Wu & Li 2019; Wu et al. 2020; Mombarg et al. 2020; Pedersen et al. 2021; Pedersen 2022b; Szewczuk et al. 2022). This can explain why many stars occur outside of the particular strips chosen to illustrate the comparison in Fig. 10. A similar conclusion was reached from the summary by Balona (2023) based on assembled space photometry over the past decade. Proper comparisons between the observed and theoretically predicted position in the HRD can help to improve the input physics and delineate free parameters used in stellar models. This is a question of various choices for opacity tables and chemical mixtures, for the mixing length parameter of (time-dependent) convection calculations, and for parameters used to describe transport processes caused by rotation, magnetic fields, and atomic diffusion, including radiative levitation (Townsend 2005; Dupret et al. 2005c, a, b; Moravveji 2016; Szewczuk & Daszyńska-Daszkiewicz 2017; Rehm et al. 2024, to list just a few papers treating some of these phenomena).
The more than 15,000 new candidate g-mode pulsators discovered from Gaia DR3 shown in Fig. 10 have global properties fully in line with the Kepler pulsators in these classes (Aerts et al. 2023), including their chemical composition (Gebruers et al. 2021; de Laverny et al. 2024). This offers great potential for ‘industrialised’ ensemble asteroseismology of thousands of intermediate- and high-mass stars. The requirement to embark upon this is that sufficient non-radial modes can be identified per star, for example from dedicated space photometry by Kepler, TESS, or PLATO, in addition to the identification of their dominant mode(s) discovered in the Gaia photometry. For now, identifications of a sufficient number of modes are only available for about 700 moderate and fast rotators observed by Kepler (Pedersen et al. 2019; Antoci et al. 2019; Cunha et al. 2019; Burssens et al. 2020; Skarka et al. 2022; Garcia et al. 2022b).
Hey & Aerts (2024) provided a path forward to extend this Kepler sample by scrutinising the light curves extracted from TESS FFIs for the 60,000 candidate pulsators in Fig. 10 with such data available. They confirmed the large majority of these 60,000 stars to be multi-periodic non-radial pulsators, and found the same dominant frequency in the TESS FFI light curves as found in the sparse Gaia DR3 light curves, highlighting Gaia’s capabilities to detect good asteroseismology targets.
Mombarg et al. (2024b) followed up on Hey & Aerts (2024) and determined the (convective core) masses, radii, and evolutionary stages for more than 10,000 g-mode pulsators from their Gaia DR3 data, after having confirmed their dominant oscillation mode from TESS FFI data. To achieve this they relied on grids of models calibrated by Kepler asteroseismology in terms of element and angular momentum transport computed by Mombarg et al. (2024a). These Gaia-based results deliver many non-radial pulsators with a mass between 2 and 3 M⊙, while none are expected from current excitation theory. This discrepancy between theory and observations will trigger further research, given the poorly known uncertainties of the Gaia DR3 observables of hot stars. Moreover, following Fig. 2, a large fraction of the stars inside and between the instability strips shown in Fig. 10 could be as-yet-unknown binaries or mergers. Their deduced luminosity and effective temperature may thus be particularly prone to systematic uncertainty as the Gaia processing ignored their potential (former) multiplicity.